OpenStudy (lia22):

A recipe that makes 2 loafs of whole wheat bread requires 1/3 cup of brown sugar. If Megan wants to make 7 loaves of bread, how much brown sugar will she require? a. 1 1/3 b. 1 1/6 c. 1 1/2 d. 1 2/3

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

$$\large \dfrac2{1/3} = \dfrac7{?}$$

jhonyy9 (jhonyy9):

for 2 loafs requires 1/3 cup for 7 ------------ x ------------------------- x = 7(1/3) /2 = (7/3)/2 = (7/3)*(1/2) = 7/6 hope this will help you

OpenStudy (skullpatrol):

2 loafs is to 1/3 cup ( as 7 loafs is to ? ) 2 divided by 1/3 = 2•3 = 6 $$\large\dfrac{2}{1/3} = \dfrac{7}{?}$$ $$\large 6 = \dfrac{7}{?}$$